if f(x)=X^2 + 8x + 1 and g(x)=2x, find the following composition
(f g)(3)

Answer by amarjeeth123(570)

fog(x)=(2x)^2+8(2x)+1=4x^2+16x+1
fog(3)=4(9)+16(3)+1=36+48+1=85